# boxedasideAside

### v1

@@ 2.1 The boundary $\overline{M}_{E,g}-M_{E,g}$
      $(E^{\prime},q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(H,q_{1}^{\prime},q_{2}^{\prime},r_{1},\cdots,r_{2g_{2}+2})$. In this case the curve has toric rank $1$ and is in $\xi_{1,g_{2}}$.
  • $\cdot$
    Case 2: $g_{1}\geq 2$. Then $\widetilde{M}_{E,g_{1};1}$ is connected. Let $(D,q_{1},q_{2})$ be a generic curve of $\widetilde{M}_{E,g_{1};1}$ and $(H,q_{1}^{\prime},q_{2}^{\prime})$ be a generic curve of ~$\widetilde{{\mathcal{H}}}_{g_{2},1}$. Then a generic curve of $\Xi_{g_{1},g_{2}}$ is of the form $(D,q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(H,q_{1},q_{2})$. Therefore, a curve in $\Xi_{g_{1},g_{2}}$ has toric rank at least $1$.
@@
Given $g_{1}\geq 2$, $g_{2}\geq 1$, we have the clutching map:
$\begin{split}\kappa^{{\rm r}}_{g_{1},g_{2}}:&\widetilde{M}_{E,g_{1}}\times\widetilde{{\mathcal{H}}}_{g_{2}}\to\widetilde{M}_{E,g_{1}+g_{2}}\\ \text{source curve: }&(D,r_{i})\times(H,r^{\prime}_{j})\mapsto(D,r_{i})\cup_{r_{2g_{1}-2}=r^{\prime}_{1}}(H,r^{\prime}_{j})\\ \text{target curve: }&(E,R_{i})\times(\mathbb{P}^{1},R^{\prime}_{j})\mapsto(E,R_{i})\cup_{R_{2g_{1}-2}=R^{\prime}_{1}}(\mathbb{P}^{1},R^{\prime}_{j})\end{split}$
where on the source curves the $(2g_{1}-2)^{th}$ smooth ramification point of $D$ is identified with the first smooth ramification point of $H$ as a node, which is still a ramification point, and similarly for the target curves. Denote the reduced image of $\kappa_{g_{1},g_{2}}^{{\rm r}}$ as $\Delta_{g_{1},g_{2}}$.
@@
The boundary of $\widetilde{M}_{E,g}$ is the union of the image of the above clutching maps. In particular,
@@
Suppose $\pi:D\to C$ is in $\overline{M}_{E,g}-M_{E,g}$. Then it must be in the image of $\widetilde{M}_{E,g}-\widetilde{M}^{\circ}_{E,g}$. Consider the forgetful morphism $\widetilde{M}_{E,g}\to\overline{\mathcal{M}}_{1,2g-2}$ that records the target curve with the $(2g-2)$ smooth branch points. Since $(C,R_{1},\ldots,R_{2g-2})$ is not smooth, $(C,R_{1},\ldots,R_{2g-2})$ is represented by a point in ¯$\overline{\mathcal{M}}_{1,2g-2}-\mathcal{M}_{1,2g-2}$, with one of its irreducible components being $E$. We know that the boundary of ¯$\overline{\mathcal{M}}_{1,2g-2}$ lies in the union of the image of the following clutching morphisms:
$\kappa_{i,j}:\overline{\mathcal{M}}_{1,i+1}\times\overline{\mathcal{M}}_{0,j+1}\to\overline{\mathcal{M}}_{1,i+j}=\overline{\mathcal{M}}_{1,2g-2},$
@@
Here $(i,j)$ runs through the pairs $i\geq 0$, $j\geq 2$, $i+j=2g-2$. We denote the reduced image of $\kappa_{i,j}$ by $\Pi_{i,j}$ and the image of $\kappa_{0}$ as $\Pi_{0}$.
Moreover, notice that $(C,R_{1},\ldots,R_{2g-2})$ can not be in $\Pi_{0}$ since $C$ should have $E$ as one of its irreducible components. Hence, there must exist $i\geq 0,j\geq 2$ such that $i+j=2g-2$ and $C=(C_{1},R_{1}^{\prime},\ldots,R_{i}^{\prime},R_{i+1}^{\prime})\cup_{R_{i+1}^{\prime}=R_{j+1}^{\prime\prime}}(C_{2},R_{1}^{\prime\prime},\ldots,R_{j}^{\prime\prime},R_{j+1}^{\prime\prime})$ for some $(i+1)$- and $(j+1)$-pointed stable curves $(C_{1},R_{1}^{\prime},\ldots,R_{i}^{\prime},R_{i+1}^{\prime})$ and $(C_{2},R_{1}^{\prime\prime},\ldots,R_{j}^{\prime\prime},R_{j+1}^{\prime\prime})$ of genera $1$ and 00, respectively. We can further choose $C_{1}$ and $C_{2}$ such that $C_{2}$ is a smooth genus 00 curve, since $C_{1}$ is allowed to be degenerate. Let $D_{1}=\pi^{-1}(C_{1})$ and $D_{2}=\pi^{-1}(C_{2})$ be the preimages of $C_{1}$ and $C_{2}$ under $\pi:D\to C$, and let $Q$ be the node on $C$ resulting from clutching. Note that $C_{2}$ has $j+1$ marked points, one of them is $Q$. We have the following possibilities for the configuration of $\pi:D\to C$.
  • $\cdot$
    Case 1: $j$ is even. Since $C_{2}$ is smooth of genus 00, any $\mathbb{Z}/2\mathbb{Z}$ cover of $C_{2}$ must have an even number of the branching points. Therefore $Q$ is not a branching point. In this case, $\pi:D_{2}\to C_{2}$ is a stable $\mathbb{Z}/2\mathbb{Z}$ cover branched at $j$ points, so this gives an element of ~$\widetilde{{\mathcal{H}}}_{\frac{j}{2}-1;1}$. Moreover, $E$ is a component of $C_{1}$ and $D_{1}\to C_{1}$ is a stable $\mathbb{Z}/2\mathbb{Z}$ cover, branched at $i$ points, hence this gives an element of $\widetilde{M}_{E,\frac{i}{2}+1;1}$. Therefore, $D\to C$ is in $\Xi_{\frac{i}{2}+1,\frac{j}{2}-1}$ and $2g-2=i+j$. The stability of $D_{2}$ implies that $j\geq 2$, so $\frac{j}{2}-1\geq 0$.
  • $\cdot$
    Case 2: $j$ is odd. Then $Q$ must be a branching point. In this case, $\pi:D_{2}\to C_{2}$ is a stable $\mathbb{Z}/2\mathbb{Z}$ cover branched at $j+1$ points, so it is in ~$\widetilde{{\mathcal{H}}}_{\frac{j+1}{2}-1}$. Moreover, $E$ is a component of $C_{1}$ and $D_{1}\to C_{1}$ is a stable $\mathbb{Z}/2\mathbb{Z}$ cover branched at $i+1$ points, hence it is in $\widetilde{M}_{E,\frac{i+1}{2}+1}$. Therefore, $D\to C$ is in $\Delta_{\frac{i+1}{2}+1,\frac{j+1}{2}-1}$ and $2g-2=i+j$. Since $i$ is odd, we have $i\geq 1$, so we have that $\frac{i+1}{2}+1\geq 2$. The stability of $C_{2}$ implies that $j\geq 2$, and it is odd so $j\geq 3$, hence $\frac{j+1}{2}-1\geq 1$.
@@
By applying the automorphism on $\mathbb{P}^{1}$, we can assume that $(Q^{\prime},R_{1},R_{2})=(\infty,0,1)$. Then, the target curve is the clutch of $(E,Q)$ and $(\mathbb{P}^{1},\infty,0,1)$ with $Q\in E$ and $\infty\in\mathbb{P}^{1}$ identified. This double cover is branched at 00 and $1$. There are two kinds of curves in $\overline{M}_{E,2}-{M}_{E,2}$:
@@
    $D=(E^{\prime},q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(\mathbb{P}^{1},q_{1}^{\prime},q_{2}^{\prime},r_{1},r_{2})$, where $E^{\prime}$ is a curve of genus $1$, such that $\pi:E^{\prime}\to E$ is an étale double cover, $\pi^{-1}(Q)=\{q_{1},q_{2}\}$, and $\pi^{\prime}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is the double cover branched at 00 and $1$, with $\pi^{\prime-1}(\infty)=\{q_{1}^{\prime},q_{2}^{\prime}\}$, $\pi^{\prime-1}(0)=r_{1}$, and $\pi^{\prime-1}(1)=r_{2}$. Here $D$ has toric rank $1$.
@@
In this example, we see from the explicit descriptions of the boundary components that they have codimension $1$ in $\overline{M}_{E,g}$. The following corollary shows that this happens for general $g$.
@@
For $g\geq 2$, the boundary of $\overline{M}_{E,g}$ has codimension $1$ in $\overline{M}_{E,g}$. Consequently, $M_{E,g}$ is dense in $\overline{M}_{E,g}$.
@@
Therefore, all loci $\Xi_{g_{1},g_{2}-1}$ and $\Delta_{g_{1},g_{2}}$ have codimension $1$ in $\overline{M}_{E,g}$, and Lemma 2.4 implies that the boundary of $\overline{M}_{E,g}$ has codimension $1$ in $\overline{M}_{E,g}$.
Suppose that $M_{E,g}$ is not dense in $\overline{M}_{E,g}$. Then there is a non-empty, open set $U\subseteq\overline{M}_{E,g}$ that does not intersect $M_{E,g}$ so that $U\subseteq\overline{M}_{E,g}-M_{E,g}$, and hence $\dim(U)\leq 2g-3$. On the other hand, Theorem A.10 says that $\overline{M}_{E,g}$ is irreducible of dimension $22g-$ 2. This implies $\dim(U)=2g-2$, which is a contradiction. To conclude, $M_{E,g}$ is dense in $\overline{M}_{E,g}$. ∎
@@ 2.2 The reduced moduli spaces $M_{E,g;n}^{{\rm red}}\subset\overline{M}_{E,g;n}^{{\rm red}}$
In this section, we introduce a “reduced” moduli space $\overline{M}_{E,g;n}^{{\rm red}}$ as a closed substack of $\overline{M}_{E,g;n}$. To construct this space, we consider the map $\pi:\overline{M}_{E,g;n}\to E$ that sums the branch points and then define $\overline{M}_{E,g;n}^{{\rm red}}$ as the fiber of this map over the identity $O_{E}\in E$. We also denote the open substack of smooth covers $M_{E,g;n}^{{\rm red}}\coloneqq M_{E,g;n}\cap\overline{M}_{E,g;n}^{{\rm red}}$.
@@
$\overline{M}_{E,g;n}^{{\rm red}}$ is a smooth proper Deligne-Mumford stack of dimension $2g-3+n$.
@@
The claim that $\overline{M}_{E,g;n}^{{\rm red}}$ is proper and Deligne-Mumford follows from the fact that $\overline{M}_{E,g;n}$ is proper and Deligne-Mumford. To prove smoothness, we observe that all the fibers of $\pi:\overline{M}_{E,g;n}\to E$ are isomorphic: to get an isomorphism between $\pi^{-1}(O_{E})=\overline{M}_{E,g;n}^{{\rm red}}$ and $\pi^{-1}(x)$, take a point $y\in E$ such that $(2g-2)\cdot y=x$ and use translation by $y$ to get an isomorphism between $\pi^{-1}(O_{E})$ and $\pi^{-1}(x)$. Because $\overline{M}_{E,g;n}$ is smooth of dimension $2g-2+n$ (Theorem A.10), the generic fiber of $\pi$ is smooth of dimension $2g-3+n$. Because all the fibers of $\pi$ are isomorphic, $\pi^{-1}(O_{E})=\overline{M}_{E,g;n}^{{\rm red}}$ is smooth of dimension $2g-3+n$, as desired. ∎
One intuitive reason to consider $\overline{M}_{E,g;n}^{{\rm red}}$ is that we ideally would like to identify double covers of $E$ if they are the same after composing with an automorphism of $E$. Because ${\rm Aut}(E)$ is an extension of a finite constant group scheme by $E$ (where $E$ acts by translations), taking the quotient by the automorphism group of $E$ will reduce the dimension of our moduli space by 1. While $\overline{M}_{E,g;n}^{{\rm red}}$ is not quite a quotient of $\overline{M}_{E,g;n}$, it is a slice for the action of ${\rm Aut}(E)$. Therefore, $\overline{M}_{E,g;n}^{{\rm red}}$ maps finitely and surjectively to $[\overline{M}_{E,g;n}/{\rm Aut}(E)]$.
@@
[…]
The first map $\varphi$ sends a cover $D\to C$ to its source curve $D$, with the $22g-$ 2 ramification points as marked points. The second map $\pi$ forgets the marked points and stabilizes the curve. Then $\pi\circ\varphi$ is quasi-finite on the smooth locus of $\overline{M}^{{\rm red}}_{E,g}$ and generically finite on each boundary component of $\overline{M}^{{\rm red}}_{E,g}$.

### v2

@@ 2.1 The boundary $\overline{M}_{E,g}-M_{E,g}$
    Suppose $g_{1}\geq 2$. Then $\widetilde{M}_{E,g_{1};1}$ is connected and only the clutching map depicted in Figure 1.(a) can occur. Let $(D,q_{1},q_{2})$ be a generic curve of $\widetilde{M}_{E,g_{1};1}$ and $(H,q_{1}^{\prime},q_{2}^{\prime})$ be a generic curve of $\widetilde{{\mathcal{H}}}_{g_{2},1}$. Then a generic curve of $\Xi_{g_{1},g_{2}}$ is of the form $(D,q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(H,q_{1},q_{2})$. Therefore, a curve in $\Xi_{g_{1},g_{2}}$ has toric rank at least $1$.
Next, we define the clutching map in the case when the clutching point is also a ramified point of the double cover, as is depicted in Figure 1.(c).
@@
Given $g_{1}\geq 2$, $g_{2}\geq 1$, we have the following clutching map:
$\begin{split}\kappa^{{\rm r}}_{g_{1},g_{2}}:&\widetilde{M}_{E,g_{1}}\times\widetilde{{\mathcal{H}}}_{g_{2}}\to\widetilde{M}_{E,g_{1}+g_{2}}\\ \text{source curve: }&(D,r_{i})\times(H,r^{\prime}_{j})\mapsto(D,r_{i})\cup_{r_{2g_{1}-2}=r^{\prime}_{1}}(H,r^{\prime}_{j})\\ \text{target curve: }&(E,R_{i})\times(\mathbb{P}^{1},R^{\prime}_{j})\mapsto(E,R_{i})\cup_{R_{2g_{1}-2}=R^{\prime}_{1}}(\mathbb{P}^{1},R^{\prime}_{j}).\end{split}$
On the source curves the $(2g_{1}-2)^{th}$ smooth ramification point of DD is identified with the first smooth ramification point of $H$ as a node, which is still a ramification point, and similarly for the target curves. Denote the reduced image of $\kappa_{g_{1},g_{2}}^{{\rm r}}$ as $\Delta_{g_{1},g_{2}}$.
@@
For $g\geq 2$, the boundary of $\widetilde{M}_{E,g}$ is the union of the image of the above clutching maps. In particular,
@@
Suppose $\pi:D\to C$ is in $\overline{M}_{E,g}-M_{E,g}$. Then it must be in the image of $\widetilde{M}_{E,g}-\widetilde{M}^{\circ}_{E,g}$. Consider the forgetful morphism $\widetilde{M}_{E,g}\to\overline{\mathcal{M}}_{1,2g-2}$ that records the target curve with the (2g2)(2g-2) smooth branch points. Since $(C,R_{1},\ldots,R_{2g-2})$ is not smooth, it is represented by a point in $\overline{\mathcal{M}}_{1,2g-2}-\mathcal{M}_{1,2g-2}$, with one of its irreducible components being $E$. We know that the boundary of $\overline{\mathcal{M}}_{1,2g-2}$ lies in the union of the image of the following clutching morphisms:
$\kappa_{1,i;0,j}:\overline{\mathcal{M}}_{1,i+1}\times\overline{\mathcal{M}}_{0,j+1}$ $\to\overline{\mathcal{M}}_{1,i+j}=\overline{\mathcal{M}}_{1,2g-2},$
@@
Here $(i,j)$ runs through the pairs $i\geq 0$, $j\geq 2$, i+j=2g2i+j=2g-2. The map $\kappa_{1,i;0,j}$ clutches two curves together at an ordinary double point, while $\kappa_{0}$ creates a node on one curve by identifying two marked points on it.
Notice that $(C,R_{1},\ldots,R_{2g-2})$ can not be in the image of $\kappa_{0}$ since $C$ should have $E$ as one of its irreducible components. Hence, there must exist $i\geq 0,j\geq 2$ such that i+j=2g2i+j=2g-2 and $C=(C_{1},R_{1}^{\prime},\ldots,R_{i}^{\prime},R_{i+1}^{\prime})\cup_{R_{i+1}^{\prime}=R_{j+1}^{\prime\prime}}(C_{2},R_{1}^{\prime\prime},\ldots,R_{j}^{\prime\prime},R_{j+1}^{\prime\prime})$ for some $(i+1)$- and $(j+1)$-pointed stable curves $(C_{1},R_{1}^{\prime},\ldots,R_{i}^{\prime},R_{i+1}^{\prime})$ and $(C_{2},R_{1}^{\prime\prime},\ldots,R_{j}^{\prime\prime},R_{j+1}^{\prime\prime})$ of genera $1$ and 0, respectively. We can further choose $C_{1}$ and $C_{2}$ such that $C_{2}$ is a smooth genus-0 curve, since $C_{1}$ is allowed to be degenerate. Let $D_{1}=\pi^{-1}(C_{1})$ and $D_{2}=\pi^{-1}(C_{2})$ be the preimages of $C_{1}$ and $C_{2}$ under $\pi:D\to C$, and let $Q$ be the node on $C$ resulting from clutching. Note that $C_{2}$ has $j+1$ marked points, one of them is $Q$. We have the following possibilities for the configuration of $\pi:D\to C$.
  1. 1.
    If $j$ is even: Write $j=2m$. Since $C_{2}$ is smooth of genus 0, any $\mathbb{Z}/2\mathbb{Z}$-cover of $C_{2}$ must have an even number of the branching points. Therefore $Q$ is not a branching point. In this case, $\pi:D_{2}\to C_{2}$ is a stable $\mathbb{Z}/2\mathbb{Z}$-cover branched at $j$ points, so this gives an element of $\widetilde{{\mathcal{H}}}_{m-1;1}$. Moreover, $E$ is a component of $C_{1}$ and $D_{1}\to C_{1}$ is a stable $\mathbb{Z}/2\mathbb{Z}$-cover, branched at i=2g2j=2(gm1)i=2g-2-j=2(g-m-1) points, hence this gives an element of $\widetilde{M}_{E,g-m;1}$. Therefore, $D\to C$ is in $\Xi_{g-m,m-1}$. The stability of $D_{2}$ implies that $j\geq 2$, so $m-1\geq 0$.
  2. 2.
    If $j$ is odd: Write $j=2m+1$. Then $Q$ must be a branching point. In this case, $\pi:D_{2}\to C_{2}$ is a stable $\mathbb{Z}/2\mathbb{Z}$-cover branched at $j+1=2(m+1)$ points, so it is in $\widetilde{{\mathcal{H}}}_{m}$. Moreover, $E$ is a component of $C_{1}$ and $D_{1}\to C_{1}$ is a stable $\mathbb{Z}/2\mathbb{Z}$-cover branched at i+1=2(gm1)i+1=2(g-m-1) points, hence it is in $\widetilde{M}_{E,g-m}$. Therefore, $D\to C$ is in $\Delta_{g-m,m}$. Since $i$ is odd, we have $i\geq 1$, implying $g-m\geq 2$. The stability of $C_{2}$ implies that $j\geq 2$, and it is odd, so $j\geq 3$. Thus we conclude $m\geq 1$.
@@
After applying an automorphism of $\mathbb{P}^{1}$, we can assume $(Q^{\prime},R_{1},R_{2})=(\infty,0,1)$. Then, the target curve is the clutch of $(E,Q)$ and $(\mathbb{P}^{1},\infty,0,1)$ with $Q\in E$ and $\infty\in\mathbb{P}^{1}$ identified. This double cover is branched at 0 and $1$. There are two kinds of curves in $\overline{M}_{E,2}-{M}_{E,2}$:
@@
    $D=(E^{\prime},q_{1},q_{2})\cup_{q_{1}=q_{1}^{\prime},q_{2}=q_{2}^{\prime}}(\mathbb{P}^{1},q_{1}^{\prime},q_{2}^{\prime},r_{1},r_{2})$, where $E^{\prime}$ is a curve of genus $1$, such that $\pi:E^{\prime}\to E$ is an étale double cover, $\pi^{-1}(Q)=\{q_{1},q_{2}\}$, and $\pi^{\prime}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is the double cover branched at 0 and $1$, with $\pi^{\prime-1}(\infty)=\{q_{1}^{\prime},q_{2}^{\prime}\}$, $\pi^{\prime-1}(0)=r_{1}$, and $\pi^{\prime-1}(1)=r_{2}$. Here $D$ has toric rank $1$.
@@
In the preceding example, the explicit descriptions of the boundary components show that they have codimension 11 in $\overline{M}_{E,g}$ when $g=2$. The following corollary demonstrates that the same holds for general $g$.
@@
For $g\geq 2$, the boundary of $\overline{M}_{E,g}$ has codimension 11 in $\overline{M}_{E,g}$. Consequently, $M_{E,g}$ is open and dense in $\overline{M}_{E,g}$.
@@
Therefore, all loci $\Xi_{g_{1},g_{2}-1}$ and $\Delta_{g_{1},g_{2}}$ have codimension 11 in $\overline{M}_{E,g}$, and Lemma 2.4 implies that the boundary of $\overline{M}_{E,g}$ has codimension 11 in $\overline{M}_{E,g}$.
Suppose that $M_{E,g}$ is not dense in $\overline{M}_{E,g}$. Then there is a non-empty, open set $U\subseteq\overline{M}_{E,g}$ that does not intersect $M_{E,g}$ so that $U\subseteq\overline{M}_{E,g}-M_{E,g}$, and hence $\dim(U)\leq 2g-3$. On the other hand, Theorem A.11 says that $\overline{M}_{E,g}$ is irreducible of dimension 2g22g-2. This implies $\dim(U)=2g-2$, which is a contradiction. To conclude, $M_{E,g}$ is dense in $\overline{M}_{E,g}$. ∎
@@ 2.2 The reduced moduli spaces $M_{E,g;n}^{{\rm red}}\subset\overline{M}_{E,g;n}^{{\rm red}}$
In this section, we introduce a “reduced” moduli space $\overline{M}_{E,g;n}^{{\rm red}}$ as a closed substack of $\overline{M}_{E,g;n}$. To construct this space, we consider the map $\pi:\overline{M}_{E,g;n}\to E$ that sends a cover $D\to C$ to the sum of the branch points of the cover. Since one component of $C$ is identified with $E$ and all other components are copies of $\mathbb{P}^{1}$, this sum of points indeed lives in ${\rm Jac}(C)\cong{\rm Jac}(E)\cong E$.
Definition 2.7.
We define $\overline{M}_{E,g;n}^{{\rm red}}$ as the fiber of the map $\pi$ over the identity $O_{E}\in E$. We also denote the open substack of smooth covers $M_{E,g;n}^{{\rm red}}\coloneqq M_{E,g;n}\cap\overline{M}_{E,g;n}^{{\rm red}}$.
@@
$\overline{M}_{E,g;n}^{{\rm red}}$ is a proper Deligne-Mumford stack of dimension 2g3+n2g-3+n.
@@
The claim that $\overline{M}_{E,g;n}^{{\rm red}}$ is proper and Deligne-Mumford follows from the fact that $\overline{M}_{E,g;n}$ is proper and Deligne-Mumford. To compute its dimension, we observe that all the fibers of $\pi:\overline{M}_{E,g;n}\to E$ are isomorphic. To get an isomorphism between $\pi^{-1}(O_{E})=\overline{M}_{E,g;n}^{{\rm red}}$ and $\pi^{-1}(x)$, take a point $y\in E$ such that $(2g-2)\cdot y=x$ and use translation by $y$ to get an isomorphism between $\pi^{-1}(O_{E})$ and $\pi^{-1}(x)$. Because $\overline{M}_{E,g;n}$ has dimension 2g2+n2g-2+n (Theorem A.11) and all the fibers of $\pi$ are isomorphic, it follows that $\pi^{-1}(O_{E})=\overline{M}_{E,g;n}^{{\rm red}}$ has dimension 2g3+n2g-3+n, as desired. ∎
One intuitive reason to consider $\overline{M}_{E,g;n}^{{\rm red}}$ is that we ideally would like to identify double covers of $E$ if they are the same after composing with an automorphism of $E$. Because ${\rm Aut}(E)$ is an extension of a finite constant group scheme by $E$ (where $E$ acts by translations), taking the quotient by the automorphism group of $E$ will reduce the dimension of our moduli space by 1. While $\overline{M}_{E,g;n}^{{\rm red}}$ is not quite a quotient of $\overline{M}_{E,g;n}$, it is a slice for the action of ${\rm Aut}(E)$. Therefore, $\overline{M}_{E,g;n}^{{\rm red}}$ maps finitely and surjectively to $\overline{M}_{E,g;n}/{\rm Aut}(E)]$. The reason we use $\overline{M}_{E,g;n}^{{\rm red}}$ rather than the quotient $\overline{M}_{E,g;n}/{\rm Aut}(E)]$ is that $\overline{M}_{E,g;n}^{{\rm red}}$ is a substack of $\overline{M}_{E,g;n}$, and hence it inherits the structure of a proper Deligne-Mumford stack. Furthermore, it is more straightforward to define the morphism in Proposition 2.9 for $\overline{M}_{E,g;n}^{{\rm red}}$ than for $\overline{M}_{E,g;n}/{\rm Aut}(E)]$.
@@
$\overline{M}_{E,g}^{{\rm red}}\xrightarrow{\varphi}\overline{\mathcal{M}}_{g,2g-2}\xrightarrow{\pi}\overline{\mathcal{M}}_{g}$
The first map $\varphi$ sends a cover $D\to C$ to its source curve $D$, with the 2g22g-2 ramification points as marked points. The second map $\pi$ forgets the marked points and stabilizes the curve. Then $\pi\circ\varphi$ is quasi-finite on the smooth locus of $\overline{M}^{{\rm red}}_{E,g}$ and generically finite on each boundary component of $\overline{M}^{{\rm red}}_{E,g}$.